in reviewing his lab book, a physics student finds the following description of a collision: "a 4 kg air hockey puck with an initial speed of 6 m/s to the right collided head on with a 1 kg puck moving to the left at the same speed. after the collision both pucks traveled to the right, the 4 kg puck at 3 m/s and the 1 kg puck at 12 m/s." is momentum conserved in this description? is kinetic energy conserved in this description? could this collision actually have taken place as described?
Answer #1

Treat motion to the right as positive and motion to the left as negative

m1 = 4 Kg , m2 = 1 Kg

v1 = 6 m/s , v2 = -6 m/s

initial momentum = m1v1 + m2v2

Initial momentum = 4 * 6 + 1 * (-6) = 18 Kg.m/s

similarly,

final momentum = 4 * 3 + 1 * 12 = 24 Kg.m/s

Therefore momentum is NOT CONSERVED

We know that when momentum is not conserved, energy is also not conserved.

I don't think this is possible because in all types of collisions, momentum is conserved unless and until there is some external force.